OpenStudy (soplaytk):

Which equations support the fact that rational numbers are closed under addition? Select each correct answer. 9√+9√=29√9+9=29 13+13=2313+13=23 6√+(−6√)=06+(−6)=0 2√+2√=22√

OpenStudy (tkhunny):

There's a whole lot of not making any sense in there. Which ones are dealing with irrational numbers? Throw those out. Note: \(\sqrt{9}\) is Rational.

OpenStudy (soplaytk):

its a multi choose awnser

OpenStudy (soplaytk):

Which equations support the fact that rational numbers are closed under addition? Select each correct answer. 9√+9√=29√9+9=29 13+13=2313+13=23 6√+(−6√)=06+(−6)=0 2√+2√=22√

OpenStudy (tkhunny):

You didn't answer my question. Reprinting the same overlapping and odd information doesn't help. Throw out the ones with Irrational Numbers. What does that leave us?

OpenStudy (soplaytk):

nvm whatever

OpenStudy (tkhunny):

So, you're not going to do it? You can solve it if you will answer my question.

OpenStudy (soplaytk):

nvm